12-72
t purchased on June 30, 2020, with a cost of $65,000, salvage value of $4,500 and a useful life ol 8 years was incorrectly entered into the depreciation system as having a useful life of 18 years Required Prepare entries to correct each of the errors a, b, and c, discovered in 2021. Ignore income taxes. Beckham Co Exercise 12-72 rporation had never been audited before December 31, 2020, the current year. Before the arrival of the auditor, the controller accounts for 2020 have not been closed. The auditors discovered that an invoice dated January 2017 for $9,000 The equipment has an estimated useful life of 10 years and no estimated residual value. prepared comparative financial statements showing the results of 2019 and 2020. The Recording and Reporting an Error, Comparative n cash at the time) was debited to 2017 operating expenses, although it was for the purchase of equipment. Statements LO6 Reported income reflected on the financial statements prepared by the company (before discovery of the er- ror) were 2017, $11,000; 2018, $22,000; 2019,$30,000; and 2020, $33,000. Disregard income tax considerations and assume that Beckham uses straight-line depreciation. Required a. Determine the correct net income for the years 2017, 2018, 2019, and 2020 b. Provide the entries to record the (1) correction of the 2020 error and (2) depreciation expense for 2020 c. Show how the correction is reported on the 2020 comparative balance sheet, retained earnings statement, and income statement. Include a proper note disclosure. On January 1, 2018, Zale Company purchased a building for $400,000. The building was estimated to have a use- Exercise 12-7 ful life of 30 years and no residual value and was depreciated using the straight-line method. In 2020, the com- Recording Error pany revised the estimated total useful life to 25 years and adjusted the residual to $5,000. In addition, in 2020, the company discovered that building improvements of $6,000 made in early 2019 were incorrectly expensed as repair expense. and Changes Estimates L Required a. Provide the journal entry to record the adjustment for the error discovered in 2020. Assume that the error is material to the company Provide the journal entry in 2020 to record depreciation expense. b. Three cases are provided below concerning a plant asset currently used in operations Exercise 1 Recording
Answer #1
12-72)
a) Correct net income for
Year 2017 = 11000 + 9000 - (9000/10) = 19100
Year 2018 = 22000 - 900 = 21100
Year 2019 = 30000 - 900 = 29100
Year 2020 = 33000 - 900 = 32100
b) Journal Entries:
Date Accounts Title Debit $ Credit $
Dec 31 2018 Equipment 9000
Operating Expense 9000
(being correction of 2017 of wrong
debiting Operating exp. Instead of Equipment)
Dec 31 2018 Depreciation Exp 3600
Acc.Dep.-Equipment 3600
(being 4 years depreciation on equipment added in
the accounting year, but purchased in 2017)
c) Effect of correction on Comparative balance sheet, retained earnings statement,
and income statement, including a proper note disclosure:
Income Statement:
Expense:
Operating exp. -9000
Depreciation 3600
Retained Earnings:
Prior Total earnings of four years           = $96000
Reduction to operating expenses           = $9000
Addition to Depreciation expense         =-$3600
Total retained earnings at the end 2020=$101400
Balance Sheet :
Assets:
Equipment $9,000
Less:Acc.Dep -3600
Net Equip $5,400
Shareholder's Equity:
Retained Earnings 101400
explanatary note: An equipment was purchased in
the year 2017, but instead of booked equipment, an
operating expense of $9000 was booked. So, the error
has been corrected in the accounting year and proper
deprecition has been made.